Consider the following partial economic information for a country where the value of the Gross Domestic Product Deflator (GDPD) is the same as the Consumer Price Index (CPI) and hence generally referred to as the Price Index. The birth rate, mortality rate, and net migrate rate are such that the population growth rate is 0%.

2010:
nominal GDP ($billions): 4486.0
price index: 108
nominal wage ($): 40,000

2011:
nominal GDP ($billions): 4710
price index: 112
nominal wage ($): 40,800

What are the real GDP figures for 2010 and 2011?
Answer #1

2010

GDP deflator = (nominal GDP/real GDP)*100.

108=(4486/real gdp) *100

108/100=4486/real gdp

1.08=4486/ real gdp

Real gdp=4486/1.08

Real gdp =4154.

2011

GDP deflator =(Nominal gdp /real gdp)*100

112=(4710/real gdp) *100

112/100=4710/real gdp

1.12=4710/real gdp

Real gdp =4710/1.12

Real gdp =4205.